A ball is thrown from a field with a speed of 20m /s at an angle of 45 with the horizontal. At what distance will it hit the field again

Grade:9

1 Answers

Arun
6 years ago
Dear Akshat
 
Horizontal Range = u² sin 2\theta/g
= 20² sin(90°)/10 = 400/10
= 40 metres
 
Hence at a distance of 40 metres it will hit the ground.
 
Regards
Arun (askIITians forum expert)